AIDA
GELINA
BRIKEN
nToF
CRIB
ISOLDE
CIRCE
nTOFCapture
DESPEC
DTAS
EDI_PSA
179Ta
CARME
StellarModelling
DCF
K40
CARME
Draft saved at 00:00:00
Fields marked with
*
are required
Entry time:
Sat Apr 5 01:28:12 2025
Author
*
:
Subject
*
:
12.46 DAQ continues G22-88/R10_14 c. 20MeV 16O2+ + 4He DSSSD bias OK - attachment 1 ambient temperature 17.7 deg C ( c. 1deg C increase cf. earlier this week ) FEE64 temperatures OK - attachment 2 All system wide checks OK WR timestamp control OK - attachment 3 ADC data item stats - attachment 4 FEE64 rates c. 0Hz except 'hot' channels aida01 ( 71k ) and aida15 ( 8k ) per FEE64 Rate and Stat spectra - attachments 5-8 events ( beam? ) aida01 asic #4 and aida04 asic #4 ( c. 200k per ASIC channel ) per FEE64 1.8.L spectra - attachments 9-10 most spectra showing background and, possibly, a small amount of scattered beam? per FEE64 1.8.W spectra - 20us FSR - attachments 11-12 aida04 3*L and 4*L spectra - attachments 13-15 distribution by strip # - uniform across asic# #4, continuing across asic #3 with abrupt fall off c. channel 9 not Rutherford - more consistent halo/beam scattering? what are the current lab angles subtended by the DSSSDs? Rutherford cross section for 20MeV 16O+4He Reaction: 4He(16O,16O)4He A: 4 16 16 4 Z: 2 8 8 2 Mass(amu): 4.002603 15.994915 15.994915 4.002603 Error(keV): 0.001 0.002 0.002 0.001 Origin: Status: Projectile lab energy: 20.000 MeV CM energy: 4.003 MeV Particle 3 Particle 4 ( scattered ) ( recoil ) Lab Cross Section CM Lab/CM Lab Lab Lab Angle Lab CM Angle Ratio Energy Angle Energy (deg) (mb/sr) (deg) (MeV) (deg) (MeV) 0.00 NaN Inf 0.000 NaN 20.000 90.000 0.000 1.00 0.143028E+09 0.572301E+07 4.999 0.040 19.976 87.500 0.024 2.00 0.894079E+07 356425. 10.017 0.040 19.902 84.992 0.098 3.00 0.176669E+07 69982.6 15.072 0.040 19.780 82.464 0.220 4.00 559329. 21951.4 20.186 0.039 19.607 79.907 0.393 5.00 229335. 8887.05 25.382 0.039 19.382 77.309 0.618 6.00 110782. 4221.76 30.690 0.038 19.103 74.655 0.897 7.00 59959.1 2235.75 36.144 0.037 18.768 71.928 1.232 8.00 35299.6 1279.52 41.790 0.036 18.371 69.105 1.629 9.00 22192.2 775.052 47.692 0.035 17.907 66.154 2.093 10.00 14729.2 489.347 53.941 0.033 17.366 63.029 2.634 11.00 10261.1 317.919 60.685 0.031 16.732 59.658 3.268 12.00 7514.20 209.725 68.185 0.028 15.976 55.907 4.024 13.00 5900.47 137.756 77.018 0.023 15.035 51.491 4.965 14.00 5693.43 85.2271 89.184 0.015 13.688 45.408 6.312 Maximum lab scattering angle ( 14.49 deg ) exceeded ________________________________________________________________________________ aida01 4*L spectra - attachments 19-23 channels 0-8 - high stats structure(s) at low energy channels 9-16 - high energy peak cf. aida04 asic #4 spectra Merger, Merger stats, TapeServer - attachments 16-18 Started Run 11 with beam ON. We observe the same pattern of counts in DSSD0 (Top left detector) but not in DSSD1,2,3 (attachment 25). Moving in the left arm with the pneumatic we see a large drop in the rate on DSSD0 to essentially zero. The beam on the CRYRING monitor appears to disappear. When we move the right pneumatic IN, we observe an increase in the number of counts in the bottom-left part of the detector. The position is likely to be a mistake in the strip assignment. No increase observed in the leakage current. Will try systematically tomorrow. Beam position monitoring from the CRYRING are shown in attachments 25-27
Encoding
:
HTML
ELCode
plain
Suppress Email notification
Attachment 1:
Drop attachments here...
Draft saved at 00:00:00
ELOG V3.1.4-unknown